OpenStudy (anonymous):

Connie has saved up $15 to purchase a new CD from the local store. The sales tax in her county is 5% of the sticker price. Write an equation and solve it to determine the value of the highest priced CD Connie can purchase with her $15, including the sales tax. Round your answer to the nearest penny.

OpenStudy (anonymous):

The equation would be 15x1.05

OpenStudy (anonymous):

If you wanted to solve it, Try multiply 15 with 0.05 and add that number to 15

OpenStudy (anonymous):

If you solve that equation, you would get exactly 15.75.

OpenStudy (anonymous):

That would look something like 15 + 15(0.05)

OpenStudy (anonymous):

= 15 + 0.75 = 15.75

OpenStudy (anonymous):

the answers could be 1. x + 0.05x = 15; x = $14.29 2. x - 0.05x = 15; x = $15.79 3. 0.5x = 15; x = $30 4. 5x = 15; x = $3.00

OpenStudy (anonymous):

Mine and @KirbyLegs 's equation are both the same. His has one extra step. For my equation, the 1 represents the $15 so you don't really need to add the other one.

OpenStudy (anonymous):

Very true. Multiplying the 0.05 with the 1 will eliminate the need to add anything extra.

OpenStudy (anonymous):

Did we solve it wrong, because none of our equations matches the ones up there.

OpenStudy (anonymous):

idk one of those answers is supposedly the answer to the question

OpenStudy (anonymous):

well ,I would go with #2 because if you round it to the nearest multiple choice answer, you would get #2

OpenStudy (anonymous):

ok thank you

OpenStudy (anonymous):

medal?